# Boundary Control and Calderón type Inverse Problems in Non-local heat equation

### v1

@@ Abstract.
We examine various density results related to the solutions of the non-local heat equation at a specific time slice, focusing on two distinct models: one with homogeneous Dirichlet boundary condition and the other with singular boundary data. We explore both the qualitative and quantitative aspects of the approximations. Additionally, we address Calderón-type inverse problems for these parabolic models, where we recover the potentials by analyzing the solutions either on the boundary or at a particular time slice. In both the density results and the Calderón type inverse problems, the Pohozaev identity plays a crucial role. Finally, in the last section, we apply the Pohozaev identity to a specific elliptic eigenvalue problem and demonstrate that the eigenfunctions, when divided by an appropriate power of the distance function, can not vanish on any non-empty open subset of the boundary.
@@ 1. Introduction
Here $\Omega\subset\mathbb{R}^{N}$ is a bounded $C^{1,1}$ domain with boundary $\partial\Omega$ and $a\in(0,1)$. The functions $u_{f},u_{F}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ are the unknowns and the potential $q\in C_{c}^{\infty}\left((0,T)\times\Omega\right)$. Furthermore, for (1), we assume the initial condition $f\in C_{c}^{\infty}(\Omega)$ and for (2) we assume the singular boundary data $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$. The distance function ${d:\Omega\to\mathbb{R}_{\geq 0}}$ is the following:
@@
The quantitative analysis enables the authors to derive an estimate for the cost of approximation, as described in [RS17]. This requires the use of variational techniques, as presented in [RS17], [FCZ00], and [Lio92]. A qualitative version of similar approximation properties for the non-local elliptic model is presented in [GSU20], where the authors investigate the Calderón type problem for the fractional heat equation. A comparable result in the $C^{k}$ space is established in [DSV19], where the authors demonstrate that any $C^{k}$ function can be approximated by smooth solutions to the non-local heat equation, with the approximation property governed by the fractional component of the equation only. However, in this article, we vary either the $N$ dimensional initial data or boundary data, and obtain $\mathrm{L}^{2}$ approximation on an $N$-dimensional interior slice. In particular, we vary the initial data $f\in C_{c}^{\infty}(\Omega)$ for the equation with homogeneous Dirichlet boundary data (1), and the singular boundary data $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$ for the equation (2). We demonstrate that for a fix time slice $T>0$, the solution sets ${\{u_{f}:f\in C_{c}^{\infty}(\Omega)\}}$ and ${\{u_{F}:F\in C_{c}^{\infty}((0,T)\times\partial\Omega)\}}$ corresponding to the equations (1) and (2) respectively, are dense in $\mathrm{L}^{2}(\Omega)$. We think such result and techniques executed here do not depend on $a\in(0,1)$, therefore this could be extended to the classical heat equation i.e. $a=1$.
@@
Let $f\in C_{c}^{\infty}(\Omega)$ and let $u_{f}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ be the solution of (1). Then, for a fix time $\tilde{T}>0$, the set $\{u_{f}(\tilde{T},\cdot):f\in C_{c}^{\infty}(\Omega)\}$ is dense in $L^{2}(\Omega)$.
@@
Let $F\in C_{c}^{\infty}((0,T)\times\Omega)$ and let $u_{F}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ be the solution of (2). Then, for a fix time $\tilde{T}>0$, the set
@@
Let $a\in(0,1)$ and let $u_{1},u_{2}\in\mathrm{L}^{2}((0,T);H^{a}(\Omega))$ satisfy: for $i=1,2$
@@
Let $a\in(0,1)$ and let $u_{1},u_{2}\in\mathrm{L}^{2}((0,T);H^{a-1(2a)}(\Omega))$ satisfy: for $i=1,2$
@@ 1.1. Preliminaries: Fractional Laplacian and fractional Sobolev space
Let $0\leq a\leq 1$. The fractional Laplacian operator $(-\Delta)^{a}$ is defined over the space of Schwartz class functions $\mathcal{S}(\mathbb{R}^{N})$ as
@@ 1.3. $\mu$ transmission spaces
$\Xi^{t}_{\pm}:\dot{H}^{s}\left(\overline{\mathbb{R}^{N}_{\pm}}\right)\xrightarrow{\sim}\dot{H}^{s-t}\left(\overline{\mathbb{R}^{N}_{\pm}}\right),\ r^{+}\Xi^{t}_{\pm}e^{+}:\overline{H}^{s}\left(\overline{\mathbb{R}^{N}_{\pm}}\right)\xrightarrow{\sim}\overline{H}^{s-t}\left(\overline{\mathbb{R}^{N}_{\pm}}\right),$
$\Lambda^{t}_{\pm}:\dot{H}^{s}\left(\overline{\Omega}\right)\xrightarrow{\sim}\dot{H}^{s-t}\left(\overline{\Omega}\right),\ r^{+}\Lambda^{t}_{\pm}e^{+}:\overline{H}^{s}\left(\overline{\Omega}\right)\xrightarrow{\sim}\overline{H}^{s-t}\left(\overline{\Omega}\right).$
@@
As $f\in C_{c}^{\infty}(\Omega)$, we have $(-\Delta)^{a}f+qf\in\mathrm{L}^{2}(\Omega)$. Hence $v\in\mathrm{L}^{2}\left((0,T);H^{a(2a)}(\overline{\Omega})\right)$ [Gru23]. Furthermore, as $f\in\mathcal{E}_{a}(\overline{\Omega})$, we have
@@
With all these regularity results in hand we are going to state integration by parts formula: If ${u,v\in H^{(a-1)(s)}(\overline{\Omega})}$ with ${s>a-\frac{1}{2}}$, then
@@
Here the boundary contributions are completely local. The formula was first proved for fractional laplacian in [Aba13] and later generalized for any classical pseudodifferential operator of order $2a$ in [Gru16a], [Gru20]. We would like to note that for any $t>0$, the integration by parts formula holds for the functions $u_{f}$ and $u_{F}$, which are the solutions to (1) and (2), respectively, where $f\in\mathrm{L}^{2}(\Omega)$ and $F\in\mathrm{L}^{2}((0,T)\times\partial\Omega)$. Thanks to theorem 7 and theorem 8, the formula clearly holds for $f\in C_{c}^{\infty}(\Omega)$ and $F\in\mathrm{L}^{2}((0,T)\times\Omega)$. Let $f\in\mathrm{L}^{2}(\Omega)$, then there exists a sequence $f_{m}\in C_{c}^{\infty}(\Omega)$ such that $f_{m}\to f$ in $\mathrm{L}^{2}(\Omega)$. For all $m\in\mathbb{N}$, we have
@@
We also mention a particular integration by parts formula which we will use throughout the article. Let $w_{1},w_{2}\in H^{a}(\mathbb{R}^{N})$ with $w_{1}\equiv w_{2}\equiv 0$ in ${\mathbb{R}^{N}\setminus\Omega}$, then
@@ 2. Density Results:
$\int_{\Omega}u_{g}(-\Delta)^{a}u_{F}-\int_{\Omega}u_{F}(-\Delta)^{a}u_{g}=-\Gamma(s)\Gamma(s-1)\int_{\partial\Omega}\frac{u_{F}}{d^{s-1}}\frac{u_{g}}{d^{s}}.$
@@
$\int_{\Omega}\partial_{t}u_{g}u_{F}+u_{g}\partial_{t}u_{F}=\int_{\Omega}\partial_{t}\left(u_{g}u_{F}\right)=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{u_{F}}{d^{a-1}}\frac{u_{g}}{d^{s}}.$
@@ 4. Further result:
While we were uploading our work to arXiv, Umberto Biccari, Mahamadi Warma, and Enrique Zuazua also posted their paper titled "Boundary Observation and Control for Fractional Heat and Wave Equations." In their work, they require the fractional power ${a\in[\frac{1}{2},1]}$ to achieve null controllability.
Acknowledgment: The author sincerely expresses gratitude to Prof. Tuhin Ghosh (Harish-Chandra Research Institute, Prayagraj, Uttar Pradesh, India) for his valuable insights and feedback, which significantly contributed to improving this article.

### v2

@@ Abstract.
We examine various density results related to the solutions of the non-local heat equation at a specific time slice, focusing on two distinct models: one with homogeneous Dirichlet boundary condition and the other with singular boundary data. In both the cases, we assume the non-local exponent $a\in(\frac{1}{2},1)$. We explore both the qualitative and quantitative aspects of the approximations. Additionally, we address Calderón-type inverse problems for these parabolic models, where we recover the potentials by analyzing the solutions either on the boundary or at a particular time slice. In both the density results and the Calderón type inverse problems, the Pohozaev identity plays a crucial role. Finally, in the last section, we apply the Pohozaev identity to a specific elliptic eigenvalue problem and demonstrate that the eigenfunctions, when divided by an appropriate power of the distance function, can not vanish on any non-empty open subset of the boundary. This particular eigenvalue problem does not need any restriction on the non-local exponent.
@@ 1. Introduction
Here $\Omega\subset\mathbb{R}^{N}$ is a bounded $C^{1,1}$ domain with boundary $\partial\Omega$ and $a\in(\frac{1}{2},1)$. The functions $u_{f},u_{F}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ are the unknowns and the potential $q\in C_{c}^{\infty}\left((0,T)\times\Omega\right)$. Furthermore, for (1), we assume the initial condition $f\in C_{c}^{\infty}(\Omega)$ and for (2) we assume the singular boundary data $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$. The distance function ${d:\Omega\to\mathbb{R}_{\geq 0}}$ is the following:
@@
The quantitative analysis enables the authors to derive an estimate for the cost of approximation, as described in [RS17]. This requires the use of variational techniques, as presented in [RS17], [FCZ00], and [Lio92]. A qualitative version of similar approximation properties for the non-local elliptic model is presented in [GSU20], where the authors investigate the Calderón type problem for the fractional heat equation. A comparable result in the $C^{k}$ space is established in [DSV19], where the authors demonstrate that any $C^{k}$ function can be approximated by smooth solutions to the non-local heat equation, with the approximation property governed by the fractional component of the equation only. However, in this article, we vary either the $N$ dimensional initial data or boundary data, and obtain $\mathrm{L}^{2}$ approximation on an $N$-dimensional interior slice. In particular, we vary the initial data $f\in C_{c}^{\infty}(\Omega)$ for the equation with homogeneous Dirichlet boundary data (1), and the singular boundary data $F\in C_{c}^{\infty}((0,T)\times\partial\Omega)$ for the equation (2). We demonstrate that for a fix time slice $T>0$, the solution sets ${\{u_{f}:f\in C_{c}^{\infty}(\Omega)\}}$ and ${\{u_{F}:F\in C_{c}^{\infty}((0,T)\times\partial\Omega)\}}$ corresponding to the equations (1) and (2) respectively, are dense in $\mathrm{L}^{2}(\Omega)$. The application of the integration by parts formula (in the $\mathrm{L}^{2}$ sense) imposes the following constraint $a\in(\frac{1}{2},1)$ on the non-local exponent. We think such result and techniques executed here could be extended to the classical heat equation i.e. $a=1$.
@@
Let $a\in(\frac{1}{2},1)$ and let $f\in C_{c}^{\infty}(\Omega)$ and let $u_{f}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ be the solution of (1). Then, for a fix time $\tilde{T}>0$, the set $\{u_{f}(\tilde{T},\cdot):f\in C_{c}^{\infty}(\Omega)\}$ is dense in $L^{2}(\Omega)$.
@@
Let $a\in(\frac{1}{2},1)$ and let $F\in C_{c}^{\infty}((0,T)\times\Omega)$ and let $u_{F}\in\mathrm{L}^{2}\left((0,T)\times\Omega\right)$ be the solution of (2). Then, for a fix time $\tilde{T}>0$, the set
@@
Let $a\in(\frac{1}{2},1)$ and let $u_{1},u_{2}\in\mathrm{L}^{2}((0,T);H^{a}(\Omega))$ satisfy: for $i=1,2$
@@
Let $a\in(\frac{1}{2},1)$ and let $u_{1},u_{2}\in\mathrm{L}^{2}((0,T);H^{a-1(2a)}(\Omega))$ satisfy: for $i=1,2$
@@ 1.1. Preliminaries: Fractional Laplacian and fractional Sobolev space
Let $10<a<$ 1. The fractional Laplacian operator $(-\Delta)^{a}$ is defined over the space of Schwartz class functions $\mathcal{S}(\mathbb{R}^{N})$ as
@@ 1.3. $\mu$ transmission spaces
$\Xi^{t}_{\pm}:\dot{H}^{s}\left(\overline{\mathbb{R}^{N}_{\pm}}\right)\xrightarrow{\sim}\dot{H}^{s-t}\left(\overline{\mathbb{R}^{N}_{\pm}}\right),\ r^{+}\Xi^{t}_{\pm}e^{+}:\overline{H}^{s}\left({\mathbb{R}^{N}_{\pm}}\right)\xrightarrow{\sim}\overline{H}^{s-t}\left({\mathbb{R}^{N}_{\pm}}\right),$
$\Lambda^{t}_{\pm}:\dot{H}^{s}\left(\overline{\Omega}\right)\xrightarrow{\sim}\dot{H}^{s-t}\left(\overline{\Omega}\right),\ r^{+}\Lambda^{t}_{\pm}e^{+}:\overline{H}^{s}\left({\Omega}\right)\xrightarrow{\sim}\overline{H}^{s-t}\left({\Omega}\right).$
@@
As $f\in C_{c}^{\infty}(\Omega)$, we have $(-\Delta)^{a}f+qf\in\mathrm{L}^{2}((0,T)\times\Omega)$. Hence $v\in\mathrm{L}^{2}\left((0,T);H^{a(2a)}(\overline{\Omega})\right)$ [Gru23]. Furthermore, as $f\in\mathcal{E}_{a}(\overline{\Omega})$, we have
@@
With all these regularity results in hand we are going to state integration by parts formula: If ${u,v\in H^{(a-1)(s)}(\overline{\Omega})}$ with ${s>a+\frac{1}{2}}$, then
@@
Here the boundary contributions are completely local. The formula was first proved for fractional laplacian in [Aba13] and later generalized for any classical pseudodifferential operator of order $2a$ in [Gru16a], [Gru20]. In articles [Gru16a], [Gru20] author shows that the boundary terms ${\partial_{\nu}\left(\frac{u}{d^{a-1}}\right),\partial_{\nu}\left(\frac{v}{d^{a-1}}\right)\in H^{s-a-\frac{1}{2}}(\partial\Omega)}$ (in case of (14) ${\frac{u}{d^{a}}\in H^{s-a-\frac{1}{2}}(\partial\Omega)}$ )))). Hence if $s>a+\frac{1}{2}$, then the $\mathrm{L}^{2}$ integral at the boundary makes sense. We would like to note that for any $a\in(\frac{1}{2},1)$ and for any $t>0$, the integration by parts formula holds for the functions $u_{f}$ and $u_{F}$, which are the solutions to (1) and (2), respectively, where $f\in\mathrm{L}^{2}(\Omega)$ and $F\in\mathrm{L}^{2}((0,T)\times\partial\Omega)$. Thanks to theorem 7 and theorem 8, the formula clearly holds for $f\in C_{c}^{\infty}(\Omega)$ and $F\in\mathrm{L}^{2}((0,T)\times\Omega)$. Let $f\in\mathrm{L}^{2}(\Omega)$, then there exists a sequence $f_{m}\in C_{c}^{\infty}(\Omega)$ such that $f_{m}\to f$ in $\mathrm{L}^{2}(\Omega)$. For all $m\in\mathbb{N}$, we have
@@
We also mention a particular integration by parts formula which we will use throughout the article. Let $a\in(0,1)$ and $w_{1},w_{2}\in H^{a}(\mathbb{R}^{N})$ with $w_{1}\equiv w_{2}\equiv 0$ in ${\mathbb{R}^{N}\setminus\Omega}$, then
@@ 2. Density Results:
$\int_{\Omega}u_{g}(-\Delta)^{a}u_{F}-\int_{\Omega}u_{F}(-\Delta)^{a}u_{g}=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{u_{F}}{d^{a-1}}\frac{u_{g}}{d^{a}}.$
@@
$\int_{\Omega}\partial_{t}u_{g}u_{F}+u_{g}\partial_{t}u_{F}=\int_{\Omega}\partial_{t}\left(u_{g}u_{F}\right)=-\Gamma(a)\Gamma(a-1)\int_{\partial\Omega}\frac{u_{F}}{d^{a-1}}\frac{u_{g}}{d^{a}}.$
@@ 4. Further result:
While we were uploading our work to arXiv, Umberto Biccari, Mahamadi Warma, and Enrique Zuazua also posted their paper titled "Boundary Observation and Control for Fractional Heat and Wave Equations." In their work, they require the fractional power ${a\in\left(\frac{1}{2},1\right)}$ to achieve null controllability.
Acknowledgment: The author was funded by the Department of Atomic Energy ((((DAE)))), Government of India, in the form of Postdoctoral Research Fellowship.
The author sincerely expresses gratitude to Prof. Tuhin Ghosh (Harish-Chandra Research Institute, Prayagraj, Uttar Pradesh, India) for his valuable insights and feedback, which significantly contributed to improving this article.
Data Availability: Data sharing is not applicable to this article as no datasets were generated or analyzed during the current study.